Suppose 0.0100 mole of Ba(OH)2 is dissolved in enough water to give 469.5 mL of solution.
What is the OH− concentration?
What is the pOH?
What is the pH?
1
Expert's answer
2020-05-10T14:58:01-0400

Solution:

Molarity of Ba(OH)2 = Moles of Ba(OH)2 / Volume of solution

Molarity of Ba(OH)2 = (0.0100 mol) / (0.4695 L) = 0.0213 M.

Molarity of Ba(OH)2 = CM(Ba(OH)2) = 0.0213 M


Barium hydroxide is a strong base for both stages of dissociation:

Ba(OH)2(s) → Ba2+ + 2OH

According to the equation: [OH] = 2 × CM(Ba(OH)2)

[OH] = 2 × (0.0213 M) = 0.0426 M

[OH] = 0.0426 M


The pOH is calculated using the expression:

pOH = − log[OH]

pOH = − log(0.0426) = 1.37

pOH = 1.37


The pH and pOH of a water solution at 25oC are related by the following equation:

pH + pOH = 14.

Then,

pH = 14 - pOH

pH = 14 - 1.37 = 12.63

pH = 12.63


Answer:

[OH] = 0.0426 M;

pOH = 1.37;

pH = 12.63.